Fermi National Laboratory
 

Operation Summary for October 9, 2009 - October 12, 2009

 

The Integrated Luminosity for the period from 10/5/09 to 10/12/09 was 45.03 pb-1.

Friday Oct. 9

The day shift began with Operations monitoring the Tevatron (TeV) store (143.63E30) and Recycler stash (130.5E10), with the Antiproton Source stacking (10.5E10), with the Linac, Booster, and Main Injector (MI) operating normally and with the NuMI and MiniBooNE experiments taking beam.

Operations terminated store 7249 at 12:41 PM.

Switchyard Septa studies began at 1:20 PM, no beam to NuMI and PBar during the study.

At 2:14 PM, CDF and D0 were making controlled accesses to their collision halls.

Beam returned to NuMI and PBar at 3:22 PM.

CDF and D0 accesses were completed at 4:31 PM.

Operations put the Tevatron (TeV) into a dry squeeze at 5:55 PM followed by a wet squeeze.

Operations put the TeV into shot setup at 7:23 PM.

Operations established store 7253 at 8:48 PM with an AIL of 224.89E30.

Saturday Oct. 10

The midnight shift began with Operations monitoring store 7253 (131.09E30) and the stash (111.1E10), with Pbar stacking (15.5E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 7253 at 9:33 AM.

D0 began a controlled access at 9:48 AM and were out of access at 11:00 AM.

Operations put the TeV into shot setup and established store 7256 at 2:27 PM with an average initial luminosity (AIL) of 257.55E30.

At 11:50 PM , a Linac bulk trim power supply died. An expert was called in, no beam to NuMI, PBar and MiniBooNE.

Sunday Oct. 11

The midnight shift began with Operations monitoring the store and the stash.

Beam returned at 2:40 AM, Linac bulk power supply had been replaced.

Operations terminated store 7256 at 9:20 AM

Operations put the TeV into shot setup and established store 7257 at 11:08 AM with an average initial luminosity (AIL) of 266.96E30.

At 2:19 PM, Operations began experiencing problems with Booster power supplies not following their ramps. Operators took the master switch and called in an expert.

Regulators were replaced in some of the Booster power supplies, beam returned at 4:19 PM.

Monday Oct. 12

The midnight shift began with Operations monitoring store 7257 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations terminated store 7257 at 4:41 AM and put the TeV into shot setup.

Store 7258 was established at 6:21 AM with an average initial luminosity (AIL) of 282.41E30.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/12/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory